OpenStudy (elise_a18):

Which of the following is not a way to represent the solution of the inequality 8x - (5x + 4) greater than or equal to -31? A number line with a closed circle on -9 and shading to the left A number line with a closed circle on -9 and shading to the right

OpenStudy (anonymous):

\[8x-(5x+4) \ge -31\] Solve for x first than any other equation.

OpenStudy (anonymous):

*like any other

OpenStudy (anonymous):

You know it's either \[x \ge 9 \] or \[x \le -9\]

OpenStudy (elise_a18):

so b?

OpenStudy (anonymous):

*both should be -9

OpenStudy (elise_a18):

so A then

OpenStudy (anonymous):

Well, what do you get for x?

OpenStudy (elise_a18):

-9

OpenStudy (anonymous):

I'm asking if you solved it, so which symbol did you get in your answer?

OpenStudy (elise_a18):

less thn or equal to

OpenStudy (anonymous):

It should be x >= -9. 8x - (5x+4) >= -31 8x -5x-4 >= -31 3x -4 >= -31 3x >= -27 x >= -9

OpenStudy (anonymous):